Circular No. 8011

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ho IN NGC 4210 Further to IAUC 8009, T. Boles reports the discovery of another ap-

parent supernova (mag 17.0) on an unfiltered CCD image taken on Nov. 5.130 UT. The new object is located at α = 12^{h}15^{m}17^{s}.97, δ = +65^{o}58′55′′.1, which is ≈ 12′′.9 east and 12′′.2 south of the center of NGC 4210. SN 2002ho was confirmed on a CCD image taken by M. Schwartz, Cottage Grove, OR, with the 0.82-m Tenagra telescope on Nov. 7.526; Boles measured mag 16.8 and position end figures 18^{s}.01, 54′′.8 from Schwartz’ image. SN 2002ho is not present on Boles’ images from Apr. 3 (limiting mag 18.5), 13, or May 27, and it is not present on Palomar Sky Survey red and blue plates.

SUPERNOVA 2002hl IN NGC 3665 Further to IAUC 8009, G. Pignata and F. Patat, on behalf of the

ERTN, report that a full reduction of a spectrum (range 380–880 nm; resolution 1.1 nm) of SN 2002hl (IAUC 8008), obtained at Calar Alto on Nov. 5.15 UT with the 2.2-m reflector (+ CAFOS), shows that this object is a typical type-Ia supernova, ∼ 2 months after maximum light, very similar to SN 1994D at a comparable phase (Patat et al. 1996, MNRAS 278, 111).

R. Foley, W. Li, and A. V. Filippenko, UCB, report that inspection of a CCD spectrum (range 310–1000 nm) of SN 2002hl, obtained on Nov. 6 UT with the Shane 3-m reflector at Lick Observatory, shows it to be a type-Ia supernova, about a month past maximum brightness.

COMET C/2002 V1 (NEAT) Additional astrometry and the following preliminary parabolic orbital

elements appear on MPEC 2002-V31.

T = 2003 Feb. 18.593 TT  ω = 152.^{o}096
Ω = 64.094 }2000.0
q = 0.09993 AU  i = 82.013
2002TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 2 5^{h}13.^{m}03 +16^{o}14.^{′}9 1.561 2.408 140.^{o}2 15.^{o}3 16.3
7 5 04.95 +16 21.6 1.436 2.331 147.1 13.4 16.0
12 4 54.69 +16 28.8 1.319 2.253 154.4 10.9 15.6
17 4 41.91 +16 35.6 1.210 2.173 162.3  7.9 15.3
22 4 26.29 +16 41.0 1.111 2.091 170.6  4.4 14.9
2002 November 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green